equation s(in Tth second)=u+at-a/2 seems to be dimensionally incorrect.why?

Grade:11

5 Answers

Akash Kumar Dutta
10 years ago

we know s=ut + 1/2.a.t^2

putting t=1

s=u + 1/2.a.............(1)

THIS EQ. IS ALSO DIMENSIONALLY INCORRECT,according to your question.

buts its not so

the eq. s(at Tth sec.)=S(t) - s(t-1)

                              = u + at -a/2

hence we see that the value of time =(t) and (t-1) has already been put.as a result it appears that the dimensions

incorrect.its the same case as of eq.(1).

IT DOES NOT VIOLATE THE DIMENSIONAL CONCEPT.

Abhishekh kumar sharma
10 years ago

the above question is

s=distance but

acoording to question the aboveequation has a unit m/sec which is independently added so we can

say it is dimensionall wrong

siddharth chandil
10 years ago

u+at-a/2

u=a(t-1/2) how can it be correct t-1/2 isn''t possible becz time -constant term is dimensinally wrong

A. Srikar
10 years ago

s=u+a(t-1/2)  is the correct formula..

 

JAYANT ARORA
9 years ago

 Yes , this is a dimensionally incorrect formula but it is derived for our help and it do violates the law of dimensions ... but we need to accept this as it is ... bcoz this helps us in problems...